a raindrop with a mesquito inside is falling at 8.5 m/s. before it hit the mesquito the drop was going at 8.8 m/s and was 30 times massive than the mesquito. What is the mesquitos average speed in g's, during the collision
1
Expert's answer
2018-12-05T15:30:10-0500

According to the law of conservation of momentum for an elastic collision

m_1 v_1+m_2 v_2=(m_1+m_2)v

Since m_1=〖30m〗_2;

v_2=(31m_2·v-30m_2·v_1)/m_2 =31·8.5-30·8.8=-0.5 m/sec

If you take the time to collision is equal to 0.03 sec

mesquitos average speed in g's = Δv/t=9/0.03=30g